# Hashed Watermark as a Filter: Defeating Forging and Overwriting Attacks in Weight-based Neural Network Watermarking

### v1

@@ V-B3 Watermark Verification
The watermark verification process is similar to the embedding process, as depicted in Fig. 3c. Concretely, upon identifying a potentially unauthorized model, the relevant subset of model parameters is extracted and subjected to hashed watermark filtering and average pooling to derive an extracted watermark $\widetilde{\mathbf{b}}$. This extracted watermark is then compared to the model owner’s watermark $\mathbf{b}$ using the watermark detection rate, which is defined by
$\rho=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}\big{[}b_{i}=\mathcal{T}(\widetilde{b}_{i})\big{]},$ (3)
@@
where $\mathcal{T}(x)$ is a threshold function that outputs $1$ if $x>0.5$ and 00 otherwise, and $\mathbbm{1}(\psi)$ is an indicator function that returns 1 if $\psi$ is true and 0 otherwise. The unauthorized model is confirmed to belong to the model owner if both of the following conditions are satisfied: (1) The watermark detection rate $\rho$ exceeds a theoretical security boundary $\rho^{\ast}$, which will be theoretically analyzed later. (2) The watermark must correspond to the output of the hash function applied to the secret key, ensuring cryptographic consistency with the predefined hash function. The watermark verification process is outlined in Algorithm 2.
Algorithm 2 Watermark Verification in NeuralMark
Watermarked model $\mathbb{M}(\theta^{*})$, secret key $\mathbf{K}$, watermark $\mathbf{b}$, index of embedding layer $\mathbf{I}_{e}$, filter rounds $R$, and security boundary $\rho^{\ast}$.
True (Verification Success) or False (Verification Failure).
Use $\mathbf{I}_{e}$ to select a subset from $\theta^{*}$ and flatten it to create $\mathbf{w}$.
for $r=1$ to $R$ do
     Perform watermark filtering on $\mathbf{w}$ to obtain $\mathbf{w}^{(r)}$.
end for
Apply average pooling on $\mathbf{w}^{(R)}$ to yield $\widetilde{\mathbf{w}}$.
Execute sigmoid mapping on $\widetilde{\mathbf{w}}\mathbf{K}$ to produce $\widetilde{\mathbf{b}}$.
Calculate watermark detection rate $\rho$ based on Eq. (3).
if $\rho\geq\rho^{\ast}$ and $\mathcal{H}(\mathbf{K})=\mathbf{b}$ then
     return True
else
     return False
end if
V-C Theoretical Analysis
We present a theoretical analysis to determine the security boundary of NeuralMark in Proposition V.1.
Proposition V.1.
Under the assumption that the hash function produces uniformly distributed outputs [38], for a model watermarked by NeuralMark with a watermark tuple $\{\mathbf{K},\mathbf{b}\}$, where $\mathbf{b}=\mathcal{H}(\mathbf{K})$, if an adversary attempts to forge a counterfeit watermark tuple $\{\mathbf{K}^{\prime},\mathbf{b}^{\prime}\}$ such that $\mathbf{b}^{\prime}=\mathcal{H}(\mathbf{K}^{\prime})$ and $\mathbf{K}^{\prime}\neq\mathbf{K}$, then the probability of achieving a watermark detection rate of at least $\rho$ (i.e., $\geq\rho$) is upper-bounded by $\frac{1}{2^{n}}\sum_{i=0}^{n-\lceil\rho n\rceil}\binom{n}{i}$.
The proof of Proposition V.1 is provided in Appendix A. Proposition V.1 provides a theoretical benchmark for establishing the security boundary of the watermark detection rate. Specifically, with $n=256$, if the watermark detection rate $\rho\geq 88.28\%$, the probability of this occurring by forgery is less than $1/2^{128}$. This negligible probability allows us to confirm ownership with high confidence. Thus, we set $n=256$ and use $88.28\%$ as the security bound for the watermark detection rate in the experiments.
VI Experiments
In this section, we evaluate NeuralMark across a variety of datasets, architectures, and tasks.
VI-A Experimental Setup
Datasets and Architectures. We use five image classification datasets: CIFAR-10, CIFAR-100 [39], Caltech-101 [40], Caltech-256 [41], and TinyImageNet [42], as well as one text generation dataset, E2E [43]. Additionally, we utilize 11 image classification architectures, including eight Convolutional architectures: AlexNet [44], VGG-13, VGG-16 [45], GoogLeNet [46], ResNet-18, ResNet-34 [47], WideResNet-50 [48], and MobileNet-V3-L [49], as well as three Transformer architectures: ViT-B/16 [50], Swin-V2-B, and Swin-V2-S [51]. Furthermore, we adopt two text generation architectures: GPT-2-S and GPT-2-M [52].
Baselines and Metrics. We compare NeuralMark with VanillaMark [13], and two state-of-the-art weight-based methods proposed in [14] and [18] (see Section II for details). For clarity, we refer to those two methods as GreedyMark and VoteMark, respectively. Additionally, we include a comparison with a method that does not involve watermark embedding, referred to as Clean. For the image classification task, we evaluate model performance using classification accuracy, while the watermark embedding task is assessed based on the watermark detection rate. As for the text generation task, we follow [53] and evaluate model performance using BLEU, NIST, MET, ROUGE-L, and CIDEr metrics, with the watermark embedding task assessed based on the watermark detection rate.
Implementation Details. We implement NeuralMark using the PyTorch framework [54] and conduct all experiments on three NVIDIA V100 series GPUs. The specific hyper-parameters are summarized below.
  • For the GPT-2-S and GPT-2-M architectures, we use the Low-Rank Adaptation (LoRA) technique [53]. Each architecture is trained for 5 epochs with a linear learning rate scheduler, starting at $2\times 10^{-4}$. We set the warm-up steps to 500, apply a weight decay with a coefficient of 0.01, and enable bias correction in the AdamW optimizer [55]. The dimension and the scaling factor for LoRA are set to 4 and 32, respectively, with a dropout probability of $0.1$ for the LoRA layers. The batch sizes for the training and test sets are 8 and 4, respectively. Moreover, we set hyper-parameter $\lambda$ to 1 and the number of filter rounds $R$ to 10.
@@ VI-A Experimental Setup
TABLE I: Comparison of classification accuracy (%) across distinct datasets using AlexNet and ResNet-18. Watermark detection rates are omitted as they all reach 100%.
@@
TABLE II: Comparison of classification accuracy (%) on CIFAR-100 using various architectures. Watermark detection rates are omitted as they all reach 100%.
@@
TABLE III: Comparison on E2E using GPT-2-S and GPT-2-M. Watermark detection rates are omitted as they all reach 100%.
@@
VI-B Fidelity Evaluation
Question 1. Is NeuralMark capable of reliably embedding watermarks while preserving model performance across a variety of datasets, architectures, and tasks?
@@ VI-B Fidelity Evaluation
Diverse Datasets. First, we evaluate the influence of watermark embedding on the model performance across diverse datasets. Table I reports the results across five image datasets using AlexNet and ResNet-18. We observe that all methods have minimal impact on model performance while successfully embedding watermarks, indicating that NeuralMark and other methods maintain model performance across diverse datasets during watermark embedding.
@@
Various Architectures. Next, we assess the impact of NeuralMark on model performance across various architectures. LABEL:{tab:fidelity_all_architectures} lists the results of NeuralMark on the CIFAR-100 dataset using VGG-13, VGG-16, GoogLeNet, ResNet-34, WideResNet-50, MobileNet-V3-L, ViT-B/16, Swin-V2-B, and Swin-V2-S. We find that NeuralMark maintains a $100\%$ watermark detection rate across a wide range of architectures while exerting minimal impact on model performance. Those observations suggest that NeuralMark is highly generalizable across various architectures.
@@
Text Generation Tasks. Finally, we evaluate the effect of NeuralMark on the text generation tasks. Table III presents the results of NeuralMark applied to the GPT-2-S and GPT-2-M architectures on the E2E dataset. We can observe that NeuralMark achieves a $100\%$ watermark detection rate while maintaining nearly lossless model performance. Those results validate the potential of NeuralMark in safeguarding the ownership of generative models.
VI-C Robustness Evaluation
Question 2. Is NeuralMark capable of withstanding forging attacks?
We adopt the setting detailed in Section IV-A to assess the robustness of NeuralMark against forging attacks. Concretely, for VanillaMark and VoteMark, we first randomly generate a counterfeit watermark and then learn the corresponding secret key by freezing the model parameters. Since GreedyMark does not require a secret key, we utilize 10 sets of randomly forged watermarks to directly verify them using the watermarked model. For NeuralMark, due to the avalanche effect of hash functions, a method similar to GreedyMark is employed, where 10 sets of randomly forged watermarks are directly verified using the watermarked model. Table IV presents the watermark detection rates of forging attacks, and we present the following significant observations. (1) For VanillaMark and VoteMark, a pair of counterfeited secret key and watermark can be successfully learned through reverse-engineering, indicating their vulnerability to forging attacks. (2) NeuralMark and GreedyMark demonstrate robust resistance against forging attacks, which aligns with our expectations. Those results suggest that NeuralMark effectively withstands forging attacks.
@@ VI-C Robustness Evaluation
TABLE IV: Comparison of watermark detection rate (%) against forging attacks using ResNet-18.
@@
TABLE V: Comparison of resistance to overwriting attacks at various trade-off hyper-parameters ($\lambda$) and learning rates ($\eta$) using ResNet-18. Values (%) inside and outside the bracket are watermark detection rate and classification accuracy, respectively.

### v2

@@ Introduction
The watermark verification process is similar to the embedding process. Concretely, upon identifying a potentially unauthorized model, the relevant subset of model parameters is extracted and subjected to hashed watermark filtering and average pooling to derive an extracted watermark $\widetilde{\mathbf{b}}$. This extracted watermark is then compared to the model owner’s watermark $\mathbf{b}$ using the watermark detection rate, which is defined by
$\rho=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\big[b_{i}=\mathcal{T}(\widetilde{b}_{i})\big],$ (2)
@@
where $\mathcal{T}(x)$ is a threshold function that outputs $1$ if $x>0.5$ and 0 otherwise, and $\mathbf{1}(\psi)$ is an indicator function that returns 1 if $\psi$ is true and 0 otherwise. The unauthorized model is confirmed to belong to the model owner if both of the following conditions are satisfied: (1) The watermark detection rate $\rho$ exceeds a theoretical security boundary $\rho^{\ast}$, which will be theoretically analyzed later. (2) The watermark must correspond to the output of the hash function applied to the secret key, ensuring cryptographic consistency with the predefined hash function. The watermark verification process is outlined in Algorithm 2 in Appendix D.
@@
Table 1: Comparison of classification accuracy (%) across distinct datasets using AlexNet and ResNet-18. Watermark detection rates are omitted as they all reach 100%.
@@
Table 2: Comparison of classification accuracy (%) on CIFAR-100 using various architectures. Watermark detection rates are omitted as they all reach 100%.
@@
Table 3: Comparison on E2E using GPT-2-S and GPT-2-M. Watermark detection rates are omitted as they all reach 100%.
@@
Security Analysis
Security Boundary Analysis
We present a theoretical analysis to determine the security boundary of NeuralMark in Proposition 1.
Propositions 1
Under the assumption that the hash function produces uniformly distributed outputs (Bellare and Rogaway 1993), for a model watermarked by NeuralMark with a watermark tuple $\{\mathbf{K},\mathbf{b}\}$, where $\mathbf{b}=\mathcal{H}(\mathbf{K})$, if an adversary attempts to forge a counterfeit watermark tuple $\{\mathbf{K}^{\prime},\mathbf{b}^{\prime}\}$ such that $\mathbf{b}^{\prime}=\mathcal{H}(\mathbf{K}^{\prime})$ and $\mathbf{K}^{\prime}\neq\mathbf{K}$, then the probability of achieving a watermark detection rate of at least $\rho$ (i.e., $\geq\rho$) is upper-bounded by $\frac{1}{2^{n}}\sum_{i=0}^{n-\lceil\rho n\rceil}\binom{n}{i}$.
The proof of Proposition 1 is provided in Appendix B. Proposition 1 provides a theoretical benchmark for establishing the security boundary of the watermark detection rate. Specifically, with $n=256$, if the watermark detection rate $\rho\geq 88.29\%$, the probability of this occurring by forgery is less than $1/2^{128}$. This negligible probability allows us to confirm ownership with high confidence. Thus, we set $n=256$ and use $88.29\%$ as the security bound for the watermark detection rate in the experiments.
Necessity of Hashed Watermark Filter
We analyze the necessity of the hashed watermark filter by comparing it to a baseline mechanism that employs a private filter rather than a hashed watermark. While this mechanism offers resistance to overwriting attacks, it remains vulnerable to forging attacks. For example, an adversary can use a $256\times 256$ identity matrix as a secret key $\mathbf{K}$ to generate a hashed watermark $\mathbf{b}$. By selecting embedding parameters $\widehat{\mathbf{w}}$ whose signs correspond to $\mathbf{b}$ (with 0 representing a negative value and 1 a positive value), the adversary can derive a private filter that selects those parameters accordingly. This allows bypassing watermark verification, i.e., $\mathcal{T}(\delta(\widehat{\mathbf{w}}\mathbf{K}))=\mathbf{b}$ and $\mathcal{H}(\mathbf{K})=\mathbf{b}$. In contrast, the hashed watermark filter cleverly intertwines the embedding parameters with the hashed watermark, rendering it essential for defending against both forging and overwriting attacks.
Experiments
In this section, we evaluate the proposed NeuralMark.
Experimental Setup
Datasets and Architectures
We use five image classification datasets: CIFAR-10 (Krizhevsky, Hinton et al. 2009), CIFAR-100 (Krizhevsky, Hinton et al. 2009), Caltech-101 (Fei-Fei, Fergus, and Perona 2004), Caltech-256 (Griffin et al. 2007), and TinyImageNet (Le and Yang 2015), as well as one text generation dataset, E2E (Novikova, Dušek, and Rieser 2017). Additionally, we utilize 11 image classification architectures, including eight Convolutional architectures: AlexNet (Krizhevsky, Sutskever, and Hinton 2012), VGG-13, VGG-16 (Simonyan and Zisserman 2015), GoogLeNet (Szegedy et al. 2015), ResNet-18, ResNet-34 (He et al. 2016), WideResNet-50 (Zagoruyko 2016), and MobileNet-V3-L (Howard et al. 2019), as well as three Transformer architectures: ViT-B/16 (Dosovitskiy 2021), Swin-V2-B, and Swin-V2-S (Liu et al. 2022). Furthermore, we adopt two text generation architectures: GPT-2-S and GPT-2-M (Radford et al. 2019).
Baselines and Metrics
We compare NeuralMark with three popular weight-based methods presented in (Uchida et al. 2017), (Liu, Weng, and Zhu 2021), and (Li et al. 2024), referred to as VanillaMark, GreedyMark, and VoteMark, respectively (see the Related Work section in Appendix A for details). Additionally, we include a comparison with a method that does not involve watermark embedding, referred to as Clean. For the image classification task, we assess model performance using classification accuracy, while the watermark embedding task is evaluated based on the watermark detection rate. As for the text generation task, we follow (Hu et al. 2022) and evaluate model performance using BLEU, NIST, MET, ROUGE-L, and CIDEr metrics, with the watermark embedding task assessed based on the watermark detection rate. More experimental details are provided in Appendix E.
Fidelity Evaluation
Diverse Datasets
@@
First, we evaluate the influence of watermark embedding on the model performance across diverse datasets. Table 1 reports the results across five image datasets using AlexNet and ResNet-18. We observe that all methods have minimal impact on model performance while successfully embedding watermarks, indicating that NeuralMark and other methods maintain model performance across diverse datasets during watermark embedding.
Various Architectures
@@
Next, we assess the impact of NeuralMark on model performance across various architectures. LABEL:{tab:fidelity_all_architectures} lists the results of NeuralMark on the CIFAR-100 dataset using ViT-B/16, Swin-V2-B, Swin-V2-S, VGG-16, VGG-13, ResNet-34, WideResNet-50, GoogLeNet, and MobileNet-V3-L. We find that NeuralMark maintains a $100\%$ watermark detection rate across a wide range of architectures while exerting minimal impact on model performance. Those observations indicate that NeuralMark exhibits a good level of generalizability across architectures.
Text Generation Tasks
@@
Finally, we evaluate the effect of NeuralMark on the text generation tasks. Table 3 presents the results of NeuralMark applied to the GPT-2-S and GPT-2-M architectures on the E2E dataset. We can observe that NeuralMark achieves a $100\%$ watermark detection rate while maintaining nearly lossless model performance. Those results demonstrate NeuralMark’s potential and generality in ownership protection of text generative models.
@@
Table 4: Comparison of detection rate (%) of counterfeit watermarks using ResNet-18.
@@
Table 5: Comparison of resistance to overwriting attacks at various trade-off hyper-parameters ($\lambda$) and learning rates ($\eta$) using ResNet-18. Values (%) inside and outside the bracket are watermark detection rate and classification accuracy, respectively. Adversary watermarks, which are consistently detected at 100%, are omitted.
@@
Table 6: Comparison of resistance to fine-tuning attacks using ResNet-18. Values (%) inside and outside the bracket are watermark detection rate and classification accuracy, respectively.
Fine-tuning Clean NeuralMark VanillaMark GreedyMark VoteMark
Robustness Evaluation
Forging Attacks
We follow the setting described in the Threat Model section to evaluate the robustness of NeuralMark against forging attacks. Specifically, for VanillaMark and VoteMark, we randomly generate a counterfeit watermark and then attempt to learn the corresponding secret key while keeping the model parameters fixed. As for GreedyMark and NeuralMark, we directly verify 10 randomly forged watermarks using the watermarked model because GreedyMark does not require a secret key, and NeuralMark benefits from the avalanche effect of the hash function and the tight coupling between the embedding parameters and the hashed watermark, making reverse-engineering infeasible. Table 4 presents the detection rates of counterfeit watermarks, from which we draw the following observations. (1) For VanillaMark and VoteMark, a pair of counterfeited secret key and watermark can be successfully learned through reverse-engineering, indicating their vulnerability to forging attacks. (2) NeuralMark and GreedyMark demonstrate robust resistance against forging attacks, which aligns with our expectations.
Overwriting Attacks
We conduct overwriting attacks targeting the watermark embedding layers, with the number of training epochs fixed at 100 to reflect limited computational resources. The optimization is guided by the loss function $\mathcal{L}_{m}+\lambda\mathcal{L}_{e}(\widetilde{\mathbf{b}},\mathbf{b}_{a})$, where $\mathbf{b}_{a}$ denotes the adversary’s watermark. Also, we analyze the effects of two key factors: the hyperparameter $\lambda$ and the learning rate $\eta$. Here, $\lambda$ controls the strength of the watermark embedding, with larger values leading to stronger embedding, while $\eta$ primarily affects model performance.
Distinct Values of $\bm{\lambda}$. We investigate the influence of $\lambda$ in overwriting attacks. Specifically, we set $\lambda$ to 1, 10, 50, 100, and 1000, respectively. Table 5 presents the results on the CIFAR-100 to CIFAR-10 task using ResNet-18. We report only the original watermark detection rate, as the adversary’s watermark detection rate reaches $100\%$. As defined in the success criterion in the Threat Model section, the original watermark must be effectively removed for overwriting attacks to be deemed successful. Thus, the overwriting attack experiments focus solely on whether the original watermark can be successfully removed. We can summarize several insightful observations. (1) As $\lambda$ increases, the original watermark detection rate of NeuralMark remains at 100%, while those of VanillaMark, GreedyMark, and VoteMark significantly decline. In particular, when $\lambda=1000$, the embedding strength of the adversary’s watermark is 1000 times greater than that of the original watermark. At this point, the original watermark detection rates for NeuralMark, VanillaMark, GreedyMark, and VoteMark on the CIFAR-100 to CIFAR-10 task are $100\%$, $53.90\%$, $49.60\%$, and $59.37\%$, respectively. Those results indicate that NeuralMark exhibits strong robustness against overwriting attacks. (2) As $\lambda$ increases, model performance remains relatively stable. This is because overwriting attacks jointly train both the main task and the watermark embedding task, enabling the model parameters to effectively adapt to both. More results are offered in Appendix F.1.
Distinct Values of $\bm{\eta}$. We examine the impact of $\eta$ in overwriting attacks. Concretely, we set $\eta$ to 0.001, 0.005, 0.01, 0.1, and 1, respectively. Table 5 lists the results on the CIFAR-100 to CIFAR-10 task using ResNet-18. We have several important observations. (1) Larger $\eta$ values hurt model performance, implying that the adversary cannot arbitrarily increase the attack strength. (2) At $\eta=0.005$, the original watermark detection rates for VanillaMark, GreedyMark, and VoteMark drop sharply, whereas NeuralMark maintains a detection rate close to $100\%$. (3) When $\eta=0.01$, the model performance of NeuralMark on the CIFAR-100 to CIFAR-10 task decreases by $2.07\%$, but its original watermark detection rate remains above the security boundary of 88.29% defined in the Security Boundary Analysis section, while those for the other methods fall significantly. (4) For $\eta>=0.1$, although the original watermark detection rate of NeuralMark drops below the security boundary, the model performance is completely compromised, indicating that the attack is ineffective. More results are provided in Appendix F.1.
Fine-tuning Attacks
We perform fine-tuning attacks on the watermark embedding layers. During the attack, the task-specific classifier is first replaced with randomly initialized parameters, after which only the parameters of the watermark embedding layers and the classifier are updated, while all other parameters remain frozen. The optimization is guided solely by the main task loss $\mathcal{L}_{m}$. Following (Liu, Weng, and Zhu 2021), we adopt the same hyper-parameters for fine-tuning attacks as during training, except for setting the learning rate to 0.001. As shown in Table 6, we find that watermarks embedded with NeuralMark maintain a $100\%$ watermark detection rate across all fine-tuning tasks. In contrast, watermarks embedded with VanillaMark, GreedyMark, and VoteMark experience a slight reduction in detection rates across several tasks. Those results indicate that fine-tuning attacks cannot effectively remove watermarks embedded with NeuralMark. Furthermore, we conduct a fine-tuning attack by updating all model parameters, as detailed in Appendix F.2.
Pruning Attacks
We evaluate the robustness of NeuralMark against pruning attacks by randomly resetting a specified proportion of parameters in the watermark embedding layer to zero. Figure 2 presents the results of NeuralMark and VanillaMark on the CIFAR-10 dataset using AlexNet and ResNet-18, respectively. As the pruning ratio increases, NeuralMark’s performance degrades slightly, while the detection rate remains nearly 100%, indicating a good level of robustness. Additional results for all baselines across different datasets are provided in Appendix F.3.
(a) NeuralMark
(b) VanillaMark
Figure 2: Comparison of resistance to pruning attacks under various pruning ratios on CIFAR-10 using AlexNet and ResNet-18.
Analysis
Parameter Distribution
Figure 3a shows the parameter distributions learned by Clean and NeuralMark on the CIFAR-100 dataset using ResNet-18. As observed, their distributions are nearly indistinguishable, making it difficult for adversaries to detect the embedded watermarks. Additional results across various architectures are provided in Appendix F.4.
Performance Convergence
Figure 3b shows the performance convergence of Clean and NeuralMark on the CIFAR-100 dataset using ResNet-18. The two curves follow a similar trajectory and remain closely aligned, indicating that NeuralMark does not hinder model convergence. Additional results across various architectures are provided in Appendix F.5.
(a) Distribution
(b) Convergence
Figure 3: Parameter distribution and performance convergence on the CIFAR-100 dataset using ResNet-18.
Filtering Rounds
To analyze watermark filtering efficacy, we generate five counterfeit watermarks and calculate the overlap ratio between parameters filtered with those and the original watermark. As shown in Figure 4, the overlap rate decreases towards zero with more filtering rounds, indicating that watermark filtering enhances the secrecy of the watermarked parameters. Furthermore, Appendix G presents additional experiments with 6 and 8 filtering rounds to evaluate their impact on NeuralMark’s effectiveness robustness against various attacks, compared to the default setting of 4. The results show that the number of filtering rounds has a negligible effect on robustness.
Additional Analyses
The impact of the watermark embedding layers and watermark length on model performance, as well as the training efficiency, is analyzed in Appendices F.6–F.8, respectively. Those results demonstrate the flexibility, effectiveness, and efficiency of NeuralMark.
Conclusion
In this paper, we present NeuralMark, a white-box method designed to protect model ownership. At the core of NeuralMark is a hashed watermark filter, which utilizes a hash function to generate an irreversible binary watermark from a secret key, subsequently employing this watermark as a filter to select model parameters for embedding. We provide a theoretical analysis of its security boundary and highlight the necessity of employing a hashed watermark as a filter. Extensive experiments on various datasets, architectures, and tasks confirm NeuralMark’s effectiveness and robustness. In future work, we plan to investigate how the proposed hashed watermark filter can be incorporated with existing watermarking approaches to offer complementary protection against broader attack scenarios.
Appendix A A. Related Work
Weight-based Method. This kind of methods (Uchida et al. 2017; Feng and Zhang 2020; Li, Tondi, and Barni 2021; Liu, Weng, and Zhu 2021; Li et al. 2024) embeds watermarks into the model parameters of neural networks. For instance, (Uchida et al. 2017) proposes the first weight-based method, which embeds the watermark into the model parameters of an intermediate layer in the neural network. Another example is (Li, Tondi, and Barni 2021), which presents a method based on spread transform dither modulation that enhances the secrecy of the watermark. However, those two methods cannot effectively resist forging and overwriting attacks. Moreover, (Feng and Zhang 2020) utilizes the secret keys to pseudo-randomly select parameters for watermark embedding and apply spread-spectrum modulation to disperse the modulated watermark across different layers. This method effectively defends against overwriting attacks while neglecting forging attacks. Additionally, (Liu, Weng, and Zhu 2021) proposes to greedily choose important model parameters for watermark embedding without an additional secret key. Although this method is effective against forging attacks, it fails to provide strong resistance to overwriting attacks of varying strength levels. Recently, (Li et al. 2024) introduces random noises into the watermarked parameters and then employs a majority voting scheme to aggregate the verification results across multiple rounds. While this method enhances the watermark’s robustness to some extent, it remains ineffective against forging and overwriting attacks.
Passport-based Method. This group of methods (Fan, Ng, and Chan 2019; Fan et al. 2021; Zhang et al. 2020; Liu et al. 2023) integrates the watermark into the normalization layers in neural networks. Specifically, (Fan, Ng, and Chan 2019; Fan et al. 2021) propose the first passport-based method, which utilizes additional passport samples (e.g., images) to generate affine transformation parameters for the normalization layers, tightly binding them to the model performance. Subsequently, (Zhang et al. 2020) integrates a private passport-aware branch into the normalization layers, which is trained jointly with the target model and is used solely for watermark verification. Recently, (Liu et al. 2023) argues that binding the model performance is insufficient to defend against forging attacks, and thus proposes establishing a hash mapping between passport samples and watermarks.
Activation-based Method. This category of methods (Rouhani, Chen, and Koushanfar 2019; Li et al. 2021; Lim et al. 2022) incorporates watermarks into the activation maps of intermediate layers in neural networks. For instance, (Rouhani, Chen, and Koushanfar 2019) incorporates the watermark into the mean vector of activation maps generated by predetermined trigger samples. Similarly, (Li et al. 2021) directly integrates the watermark into the activation maps associated with the trigger samples. Additionally, (Lim et al. 2022) embeds the watermark into the hidden memory state of a recurrent neural network.
Appendix B B. Proof for Proposition 1
Proposition 1. Under the assumption that the hash function produces uniformly distributed outputs (Bellare and Rogaway 1993), for a model watermarked by NeuralMark with a watermark tuple $\{\mathbf{K},\mathbf{b}\}$, where $\mathbf{b}=\mathcal{H}(\mathbf{K})$, if an adversary attempts to forge a counterfeit watermark tuple $\{\mathbf{K}^{\prime},\mathbf{b}^{\prime}\}$ such that $\mathbf{b}^{\prime}=\mathcal{H}(\mathbf{K}^{\prime})$ and $\mathbf{K}^{\prime}\neq\mathbf{K}$, then the probability of achieving a watermark detection rate of at least $\rho$ (i.e., $\geq\rho$) is upper-bounded by $\frac{1}{2^{n}}\sum_{i=0}^{n-\lceil\rho n\rceil}\binom{n}{i}$.
[…]
Appendix C C. Workflow of NeuralMark
Figure 5 illustrates the workflow of NeuralMark, including watermark generation, embedding, and verification stages.
(a) Generation
(b) Embedding
(c) Verification
Appendix D D. Algorithms of NeuralMark
Algorithms 1-2 offer the watermark embedding and verification processes in NeuralMark, respectively.
Algorithm 1 Watermark Embedding in NeuralMark
1:Training dataset $\mathcal{D}$, secret key $\mathbf{K}$, index of embedding layer $\mathbf{I}_{e}$, hyper-parameters $\lambda$, $T$, and filter rounds $R$.
2:Watermarked model $\mathbb{M}(\theta^{*})$.
3:Randomly initialize the model parameter $\theta$.
4:Generate the watermark $\mathbf{b}=\mathcal{H}(\mathbf{K})$.
5:for $t=0$ to T1T-1 do
6:  Use $\mathbf{I}_{e}$ to select a subset from $\theta$ and flatten it into $\mathbf{w}$.
7:  for $r=1$ to $R$ do
8:   Perform watermark filtering on $\mathbf{w}$ to obtain $\mathbf{w}^{(r)}$.
9:  end for
10:  Apply average pooling on $\mathbf{w}^{(R)}$ to yield $\widetilde{\mathbf{w}}$.
11:  Execute sigmoid mapping on $\widetilde{\mathbf{w}}\mathbf{K}$ to produce $\widetilde{\mathbf{b}}$.
12:  Update $\theta$ based on Eq. (1).
13:end for
Algorithm 2 Watermark Verification in NeuralMark
1:Watermarked model $\mathbb{M}(\theta^{*})$, secret key $\mathbf{K}$, watermark $\mathbf{b}$, index of embedding layer $\mathbf{I}_{e}$, filter rounds $R$, and security boundary $\rho^{\ast}$.
2:True (Verification Success) or False (Verification Failure).
3:Use $\mathbf{I}_{e}$ to select a subset from $\theta^{*}$ and flatten it to create $\mathbf{w}$.
4:for $r=1$ to $R$ do
5:  Perform watermark filtering on $\mathbf{w}$ to obtain $\mathbf{w}^{(r)}$.
6:end for
7:Apply average pooling on $\mathbf{w}^{(R)}$ to yield $\widetilde{\mathbf{w}}$.
8:Execute sigmoid mapping on $\widetilde{\mathbf{w}}\mathbf{K}$ to produce $\widetilde{\mathbf{b}}$.
9:Calculate watermark detection rate $\rho$ based on Eq. (2).
10:if $\rho\geq\rho^{\ast}$ and $\mathcal{H}(\mathbf{K})=\mathbf{b}$ then
11:  return True
12:else
13:  return False
14:end if
Appendix E E. Implementation Details
We implement NeuralMark using the PyTorch framework (Paszke et al. 2019) and conduct all experiments on three NVIDIA V100 series GPUs. The specific hyper-parameters are summarized below.
  • For the GPT-2-S and GPT-2-M architectures, we use the Low-Rank Adaptation (LoRA) technique (Hu et al. 2022). Each architecture is trained for 5 epochs with a linear learning rate scheduler, starting at $2\times 10^{-4}$. We set the warm-up steps to 500, apply a weight decay with a coefficient of 0.01, and enable bias correction in the AdamW optimizer (Loshchilov, Hutter et al. 2017). The dimension and the scaling factor for LoRA are set to 4 and 32, respectively, with a dropout probability of $0.1$ for the LoRA layers. The batch sizes for the training and test sets are 8 and 4, respectively. Moreover, we set hyper-parameter $\lambda$ to 1 and the number of filter rounds $R$ to 10.
Appendix F F. Additional Experimental Results
F.1 Overwriting Attacks
Table 7 lists the results of the overwriting attack on the CIFAR-10 to CIFAR-100 task using ResNet-18, which are consistent with those observed on the CIFAR-100 to CIFAR-10 task reported in the main text. Those results further demonstrate that NeuralMark exhibits strong robustness against overwriting attacks across a range of attack strengths.
Table 7: Comparison of resistance to overwriting attacks at various trade-off hyper-parameters ($\lambda$) and learning rates ($\eta$) using ResNet-18. Values (%) inside and outside the bracket are watermark detection rate and classification accuracy, respectively.